Please help me find the volume of a shape. The shape in my book is a cone on top of a cylinder. The base of the cylinder is 8 and the height is 14. The height of the cone is 21. I've tried many different ways and cannot seem to come up with the correct answer. The book gives me an answer of 1056 cm^3
Thank you!

Answer by stanbon(75887)

The shape in my book is a cone on top of a cylinder. The base of the cylinder is 8 and the height is 14. The height of the cone is 21. I've tried many different ways and cannot seem to come up with the correct answer. The book gives me an answer of 1056 cm^3
----------------
Cylinder Volume: V = pi*r^2*h = pi*4^2*14 = 703.717
---
Cone Volume: V = (1/3)pi*r^2*h = (1/3)pi*4^2*21 = 351.858
----
Total volume = 703.717+351.858 = 1055.58 cm^3
Rounding up you get Total volume = 1056 cm^3
=================================================
Cheers,
Stan H.